# Hyperparameter Optimization in Machine Learning

### v1

@@ 3.3 Quasi-random search
The vast majority of model-based methods is, in principle, sequential: the surrogate model suggests the next point to evaluate while being simultaneously adapted online as new data from the execution of $A$ is collected and processed. We can model such iterative process introducing a two mappings $\pi_{1}$ and $\pi_{2}$, one corresponding to the iteration in the hyperparameter space, the other to the update of the surrogate model:
$\left(\begin{array}[]{c}\lambda_{k+1}\\ \hat{f}_{k+1}\end{array}\right)=\left(\begin{array}[]{c}\pi_{1}(\hat{f}_{k},\upsilon)\\ \pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\hat{f}_{k},\upsilon\right)\end{array}\right).$ (4.2)
Another possibility is to first learn $\hat{f}$ and then use it to suggest one or several hyperparameter configurations in a second stage, following a parallel scheme. Examples of this latter approach include [feurer2015initializing, fusi2018probabilistic]. The methods, however, assume the presence of multiple tasks and are closer in nature to meta or transfer learning settings. Still, even in the sequential approach, typically $\hat{f}_{0}$ is initialized by using several evaluations obtained in a “warm-up stage”, e.g., by random search or Latin hypercube sampling. While the foremost representatives of these type of methods – tied to BO – are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in Chapter 5.
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an acquisition function depending on the surrogate model, as we will see in Section 4.2. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of MB techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
@@ 4.1 Bayesian Optimization
BO has long been used as a method to globally optimize black-box functions (i.e., functions for which we only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda)$, and dates back at least to the seventies with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. While limited to small-to-medium dimensional settings, BO is an effective approach to optimize (highly) multi-modal, non-convex and possibly expensive-to-compute functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give a brief overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for an intuitive introduction of the mathematical aspects of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning.
BO is a model-based technique, whose surrogate model is defined by a stochastic process, representing a probability distribution over the objective $f$. This surrogate model is fitted to available observations by Bayesian conditioning, so that afterwards it represents the Bayesian posterior distribution over function values $f(\lambda)$ away from the observations. This posterior distribution is used to suggest the next value to probe. Typically, the suggestion rule is expressed as the maximization of an acquisition function, balancing exploration – collecting information to improve the model itself, by proposing points away from previous observations – and exploitation – sampling close to the current minimizer in order to refine the best solution. Referring to the notation of the previous section and to Equation (4.2), the first component $\pi_{1}$ constitutes the decision rule based on an acquisition function, while the second component $\pi_{2}$ maps to the Bayesian update of our surrogate model given the data observed so far. Next, we describe a “classical” instantiation of BO, whereby the surrogate model is a Gaussian process [williams2006gaussian] and the acquisition function is the expected improvement [mockus1978application]. This approach is well-suited for the global optimization of smooth multivariate real functions, and constitutes the core of the Spearmint HPO package [snoek_practical_2012].
@@ 4.1.1 Bayesian linear model
where $\varphi:\Lambda\to\mathcal{V}$ is a feature map. While using linear maps directly in the input space would result in an overly simplistic model – we are trying to model an objective, the validation error, which can be (highly) non-convex and multimodal – working in a different feature space allows us to capture much richer behaviors, but still retain most of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in (4.3) models potential errors in the observations. In HPO, this noise arises because the validation error is only an estimate of the true generalization error. Additionally, this noise may partially account for the stochastic nature of the learning algorithm itself. It is important to distinguish between aleatoric uncertainty, which derives from inherent randomness in the data, and epistemic uncertainty, which arises from lack of knowledge about the model. A heteroscedastic approach can help disentangle these types of uncertainties, as proposed by [griffiths:2021] in their work on achieving robustness to aleatoric uncertainty with heteroscedastic Bayesian optimization.
Suppose now we are at the $k$-th iteration of the algorithm, after having already collected $k$ observations. Let $Y_{k}=\left(f(\lambda_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ be the vector of observed values, and denote by $X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)$ the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
@@
where $x_{*}=\varphi(\lambda_{*})$ is the query point in the feature space. We could follow an empirical risk minimization approach, obtaining the weights $w_{k}\in\mathbb{R}^{\mathrm{dim}(\mathcal{V})}$ of the model (4.3) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k+1}=h_{w_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w_{k})$, capable of capturing only our belief of the amount of noise present in the observations. Rather, we would like the stochasticity of $\hat{f}$ to (primarily) represent our uncertainty about the behavior of $f$ in unexplored regions.111 The following argument, mutated from jones1998efficient, may help in better grasping the difference between these two aspects. Consider the case where $f$ is a non-linear smooth deterministic function. We may then either omit the noise term $\varepsilon$ in (4.3), or leave it for capturing modeling errors. In the first scenario, the ERM approach would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously sampled. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k+1}$ would banally be inaccurate at $\lambda_{i}$ for $i\leq k$. Thus, informally speaking, both ERM modeling options would leave us with a surrogate model that is surely inaccurate for virtually all $f$’s. A way to escape this apparent modeling stalemate is to replace the independent Gaussian noise with one that depends on the evaluation point, i.e. $\varepsilon=\varepsilon(\lambda)$. Then, when fitting the model to the $k$ observations, one may have $\varepsilon_{k}(\lambda_{i})=0$ for $i\leq k$ while letting $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, by continuity ($\varepsilon_{k}$ is smooth and continuous as difference of $C^{1}$ functions), $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behavior. Note, however, that the noise term $\varepsilon$ in (4.3) remains independent and indeed represents, there, the measurement noise. To this aim, we take a Bayesian perspective on the fitting problem, and reason about all the linear models (in the feature space) that agree with the observations, weighted by their “realization probability”.
@@
$\hat{f}_{k+1}(\lambda_{*})=\int p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)p_{w}(\cdot|X_{k},Y_{k})\,\mathrm{d}w.$ (4.4)
@@
Regarding the numerator of (4.6), we have that, for the same argument of (4.3),
@@
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with (4.7), we can now also compute the denominator of (4.6) by marginalizing over $w$ [shahriari:2016:taking_human_out], obtaining
$p_{y}(\cdot|X_{k})=\int p_{y}(\cdot|X_{k},w)p_{w}\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I)$ (4.8)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the corresponding kernel
@@
and rewrite (4.8) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathbb{R}^{k\times k}$ as
@@
We finally have all the terms to compute (4.4); which is, again, Gaussian [williams2006gaussian]:
$\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k})=\mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)k_{*})$ (4.10)
@@
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$ and the Matérn family, which contains kernels (defined as functions of $r_{\Sigma}$, above) with various degree of smoothness [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow to sidestep the definition of (possibly obscurer) feature maps and covariance matrices. The series of “free parameters”222 That is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel. that appears in (4.10) – usually called hyperparameters in the Bayesian literature (not to be confused with $\lambda$) – do not have to be fixed constants, but may be estimated from the observations [williams2006gaussian, shahriari:2016:taking_human_out] (automatic relevance determination). Thus, at this stage, the design choice left to the experimenter mainly boils down the choice of the kernel function.
@@ 4.1.2 Gaussian Processes
$\hat{f}=\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta).$
@@
$\hat{f}_{k+1}=\pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\xi_{2}\right)=\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta\,|\,X_{k},Y_{k}),$ (4.11)
which expresses the update as the conditioning of the Gaussian process on the observations $X_{k}$, $Y_{k}$. While this posterior is in principle an infinite-dimensional object, we only ever need its marginal distributions over single points $\lambda_{*}$ (or over small sets, such as pairs) in order to compute acquisition functions and make sampling decisions. These marginals are Gaussian with mean and (co)variance given by simple linear algebra expressions (4.10). More details on acquisition functions and the decision step $\pi_{1}$ is provided in Section 4.2.
To summarize, Equations (4.12) and (4.11) specify a particular simple instantiation of a GP-based BO algorithm that may be used for HPO. The configuration parameters $\nu_{2}$ may contain choices regarding the mean and covariance functions of the GP, the method with which the GP hyperparameters are estimated, and the acquisition function to use. Apart from the simplicity of a Gaussian process setup, its uncertainty representation fares particularly well for small number of observations, and the posterior representation is updated efficiently and robustly once additional data becomes available.
4.1.3 Limitations
@@
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions, and take on the scalability issues in snoek_scalable_2015 by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms and bergstra_making_2013 focus both on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators (see also Section 4.3). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next section) to improve performance in the first stage of HPO, when the surrogate model may still be rather unreliable.
4.2 Acquisition Functions
Suppose we fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $C=\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function (AF) $\alpha(\lambda|\hat{f}_{k})$. In other words, our suggestion mapping $\pi_{1}$ becomes

### v2

@@ 4.1 Model-based HPO: a general schema
The vast majority of model-based methods is, in principle, sequential: the surrogate model suggests the next point to evaluate while being simultaneously adapted online as new data from the execution of the learning algorithm $A$ is collected and processed.
It can be useful to think about a MB-HPO method as an iterative process over both the hyperparameter space $\Lambda$ and the surrogate function space $\hat{\mathcal{F}}$. To represent this process, we introduce two mappings $\pi_{1}:\mathcal{F}\times\mathcal{X}\to\Lambda$ and $\pi_{2}:(\Lambda\times\mathbb{R})^{*}\times\hat{\mathcal{F}}\times\mathcal{X}\to\hat{\mathcal{F}}$, one corresponding to the iteration in the hyperparameter space, the other to the update of the surrogate model. Then, we can describe the iterates of a MB-HPO method as follows:
$\left(\begin{array}[]{c}\lambda_{k+1}\\ \hat{f}_{k+1}\end{array}\right)=\left(\begin{array}[]{c}\pi_{1}(\hat{f}_{k},\upsilon_{k})\\ \pi_{2}\left(\mathcal{C}_{k},\hat{f}_{k},\upsilon_{k}\right)\end{array}\right),$ (4.3)
where the variables $\nu_{k}$ represent auxiliary inputs and $\mathcal{C}_{k}$ a set of paris of configurations and response function evaluations. The surrogate update mapping $\pi_{2}$ may often entail an online update of $\hat{f}$, therefore depending only on the last collected point $(\lambda_{k},f(\lambda_{k}))$ and the current surrogate model $\hat{f}_{k}$.
Typically $\hat{f}_{0}$ is initialized by fitting it to several evaluations obtained in a “warm-up stage”, e.g., where configuration are generated by random search or Latin hypercube sampling. Then, the points $\mathcal{C}_{k}=\mathcal{C}_{0}\cup\{(\lambda_{j},f(\lambda_{j}))\}_{j=1}^{k}$ that comprise past configurations and evaluations of the response function (including warm-up evaluations) can be though of as a dataset to fit the surrogate model. We summarize the typical steps of a model-based method in Algorithm 1, where the decision map $\pi_{i}$ is implemented by the optimization of an AF $\alpha$. The algorithm generates a sequence of $T$ configurations. As in the previous chapter, we denote by $\hat{\lambda}_{k}$ the best configuration found after $k$ steps.
@@ 4.2 Bayesian optimization
BO has long been used as a method to globally optimize black-box functions, that is, functions for which we may only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda)$. It dates back at least to the seventies with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. While limited to small-to-medium dimensional settings, BO is an effective approach to optimize (highly) multi-modal, non-convex and possibly expensive-to-compute functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give a brief overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for an intuitive introduction of the mathematical aspects of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning.
In Bayesian optimization, $\hat{y}$ is a stochastic process rather than a function and $\hat{y}(\lambda)$ is a random variable. Here randomness arises because of the epistemic uncertainty of the surrogate model (due to the small size of $\mathcal{D}$) about the response at $\lambda$. This surrogate model is fitted to available observations by Bayesian conditioning, so that afterwards it represents the Bayesian posterior distribution, which is used to decide the next value to probe. Such decision is most often expressed as the maximization of an acquisition function, as we introduced before. The AF balances exploration — collecting information to improve the model itself, by proposing points away from previous observations — and exploitation — picking points close to the current minimizer in order to refine the best solution. Thus, referring to Eq. (4.3), a typical BO method implements $\pi_{1}$ with the optimization of an AF and $\pi_{2}$ with a Bayesian update of the stochastic surrogate model.
First, we discuss standard choices to design and fit surrogate models in BO, including Gaussian processes and tree Parzen estimators. Then, we focus on AFs, especially on the expected improvement [mockus1978application]. One may say that a most classic instantiation of BO consists in modelling the surrogate with a Gaussian process and using as AF the expected improvement. This approach is well-suited for the global optimization of smooth multivariate real functions, and it provides an accessible setting to approach BO for hyperparameter optimization. Nevertheless, it may not be the most suitable for some HPO problems; thus we discuss limitations along the way.
@@ 4.2.1 Bayesian linear model
where $\varphi:\Lambda\to\mathcal{V}$ is a feature map. While using linear maps directly in the input space would result in an overly simplistic model – we are trying to model an objective, the validation error, which can be (highly) non-convex and multimodal – working in a different feature space allows us to capture much richer behaviors but still retain most of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in Eq. (4.4) models potential errors in the observations (see also 4.2.4).
Suppose now we are at the $k$-th iteration of the algorithm, after having already collected $k$ observations $\mathcal{C}=\{(\lambda_{i},f(\lambda_{i}))\}_{i\in[k]}$. Let $Y_{k}=\left(f(\lambda_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ be the vector of observed values, and denote by $X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)$ the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
@@
where $x_{*}=\varphi(\lambda_{*})$ is a query point in the feature space. This is going to be useful later in Section 4.2.3, where an AF employs $\hat{f}_{k}$ to select the next configuration $\lambda_{k+1}$.
Modelling uncertainty away from observations.
We could follow an empirical risk minimization approach, obtaining the weights $w_{k}\in\mathbb{R}^{\mathrm{dim}(\mathcal{V})}$ of the model Eq. (4.4) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k+1}=h_{w_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w_{k})$, capable of capturing only our belief of the amount of noise present in the observations. Rather, we would like the stochasticity of $\hat{f}$ to (primarily) represent our uncertainty about the behavior of $f$ in unexplored regions. To better grasp the difference between these two aspect, consider the case where $f$ is a (non-linear) smooth deterministic function.  [^1] The following argument is mutated from jones1998efficient. If we were to follow an ERM approach, we could either omit the noise term $\varepsilon$ in Eq. (4.4), or leave it for capturing modelling errors. However, the first scenario, would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously observed. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k+1}$ would trivially be inaccurate at $\lambda_{i}$ for $i\leq k$ (as $f$ is deterministic). Thus, both ERM modelling options would leave us with a surrogate model that is almost surely inaccurate.
A way to escape this apparent stalemate is to replace the independent Gaussian noise with one that depends on the evaluation point, i.e. $\varepsilon=\varepsilon(\lambda)$. Then, when fitting the model to the $k$ observations, one may have $\varepsilon_{k}(\lambda_{i})=0$ for $i\leq k$ while letting $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, by continuity ($\varepsilon_{k}$ is smooth and continuous as difference of $C^{1}$ functions), $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behaviour.
Thus, we can reason about all the linear models (in the feature space) that agree with the observations, weighted by their “realization probability”.
@@
$\hat{f}_{k+1}(\lambda_{*})=\int p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)p_{w}(w|X_{k},Y_{k})\,\mathrm{d}w.$ (4.5)
@@
Regarding the numerator of Eq. (4.7), we have that, for the same argument of Eq. (4.6),
@@
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with Eq. (4.8), we can now also compute the denominator of (4.7) by marginalizing $p_{y}(\cdot|X_{k})$ over $w$. We obtaining
$p_{y}(\cdot|X_{k})=\int p_{y}(\cdot|X_{k},w)p_{w}(w)\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I)$ (4.9)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the kernel corresponding to the feature map $\phi$ as follows:
@@
and rewrite Eq. (4.9) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathbb{R}^{k\times k}$ as
@@
We finally have all the terms to compute (4.5); which is, again, Gaussian:
$\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k})=\\ \mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)k_{*})$ (4.11)
@@
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$ and the Matérn family, which contains kernels (defined as functions of $r_{\Sigma}$, above) with various degree of smoothness [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow to sidestep the definition of (possibly obscurer) feature maps and covariance matrices. The series of “free parameters”222 That is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel. that appears in (4.11) – usually called hyperparameters in the Bayesian literature (not to be confused with $\lambda$) – do not have to be fixed constants, but may be estimated from the observations [williams2006gaussian, shahriari:2016:taking_human_out] (automatic relevance determination). Thus, the design choice left to the experimenter mainly boils down the choice of the kernel function.
@@ 4.2.2 Gaussian processes
$\hat{f}\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta).$
@@
$\hat{f}_{k+1}=\pi_{2}\left(\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k},\xi_{2}\right)\sim\mathcal{GP}(0,\kappa_{\Sigma}+\sigma_{\varepsilon}^{2}\delta\,|\,X_{k},Y_{k}),$ (4.12)
which expresses the update as the conditioning of the Gaussian process on the observations $X_{k}$, $Y_{k}$. While this posterior is in principle an infinite-dimensional object, we only ever need its marginal distributions over single points $\lambda_{*}$ (or over small sets, such as pairs) in order to compute AFs and make sampling decisions. These marginals are Gaussian with mean and (co)variance given by simple linear algebra expressions (4.11).
Limitations of Gaussian processes.
@@
snoek_practical_2012 tackle the first problem, proposing an extension of the expected improvement that accounts for pending executions, and take on the scalability issues in snoek_scalable_2015 by replacing the GP with a Bayesian neural network. hutter_sequential_2011, bergstra2011algorithms and bergstra_making_2013 focus both on scalability and applicability to discrete and conditional search spaces by using random forests and tree Parzen estimators (see also Section 4.2.5). More recent advances in the field [springenberg2016bayesian, wu2017bayesian, klein2017fast] focus on replacing GPs with more practical and scalable approaches, exploiting additional information besides function evaluations (thus “dropping” the black-box assumption for $f$). The additional information could be, for example, the performance of $A$ on a small subsets of $D$ [klein2017fast] or (possibly noisy) gradients [wu2017bayesian]. Finally, falkner2018bohb propose hybridizing BO with bandit-based methods (discussed in the next section) to improve performance in the first stage of HPO, when the surrogate model may still be rather unreliable.
4.2.3 Acquisition functions
Suppose we fitted our surrogate model $\hat{f}_{k}$ to the current set of evaluations $\mathcal{D}=\{\lambda_{j},f(\lambda_{j})\}_{j=1}^{k}$. To decide which point to evaluate next, BO optimizes an acquisition function $\alpha(\lambda|\hat{f}_{k})$. In other words, our suggestion mapping $\pi_{1}$ becomes